OpenStudy (superhelp101):

11x4+35x2+6 I want to use the method where you multiply the first and last to factor. Any help?

OpenStudy (superhelp101):

@dumbcow can you help me?

OpenStudy (superhelp101):

@tkhunny

OpenStudy (dumbcow):

ok 11*6 = 66 33+2 = 35 33*2 = 66 (11x^4 +33x^2) +(2x^2 + 6) 11x^2 (x^2 +3) + 2(x^2 +3) (x^2 +3)(11x^2 +2)

OpenStudy (superhelp101):

ok thank you!

OpenStudy (dumbcow):

yw

OpenStudy (tkhunny):

One would ALWAYS multiply first and last - and every other combination.

OpenStudy (superhelp101):

oh i see, thanks!

OpenStudy (superhelp101):

@dumbcow so I noticed that you used grouping. Would that always work?

OpenStudy (tkhunny):

"Grouping" is rarely beneficial and rarely has application in the real world. Most of the time, "grouping" is accomplished by knowing beforehand where things need to be. I do not recommend spending much time on it. It does occasionally crop up as useful, but not very often.

OpenStudy (superhelp101):

in these type of problems, like factoring. It is a bit easier for me on grouping, then using the wrong method. Do you mind showing me the wrong way on this problem. So i will get introduced. ;)